OpenStudy (anonymous):

A line passes through the points(-10,-4) and(-1,2) . What is the y-intercept of the line?

OpenStudy (the_fizicx99):

Ok, find the slope then put it into point slope form then convert that into slope intercept form, which is y = mx + b where b is the y intercept

OpenStudy (the_fizicx99):

To find the slope put your coordinates into M = (Y2-y1)/(X2-x1) M = 2-(-4)/-1-(-10) M = 2 +4/ -1+10 Solve that

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -10}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ -1}}\quad ,&{\color{blue}{ 2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green} m}(x-{\color{red}{ x_1}})\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (the_fizicx99):

=.=

OpenStudy (jdoe0001):

you'd end up with something like \(\large \begin{array}{llll} y=\square x&+\square \\ &\ \uparrow \\ &\textit{y-intercept} \end{array}\)

OpenStudy (anonymous):

so my y-intercept is -4

OpenStudy (the_fizicx99):

Well did you find your slope first?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

slope was 2/9

OpenStudy (the_fizicx99):

6/9

OpenStudy (anonymous):

6/9 was not none of the answers

OpenStudy (the_fizicx99):

Simplified is 2/3 So, y-1y=2/3(x-x1) y -(-4)=2/3x + 6.7 y + 4 = 2/3x + 6.7 -4 -4 y = 2/3x + 2.7 What're the options, I'm pretty sure 0.66666666666666666666666666666667 rounded is 6.7

OpenStudy (the_fizicx99):

Typo, 6.6666666666666666666666666666667 *

OpenStudy (the_fizicx99):

What're the options, I wanna make sure :p

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -10}}\quad ,&{\color{blue}{ -4}})\quad &({\color{red}{ -1}}\quad ,&{\color{blue}{ 2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ 2}}-({\color{blue}{ -4}})}{{\color{red}{ -1}}-({\color{red}{ -10}})}\implies \cfrac{6}{9} \\ \quad \\ y-({\color{blue}{ -4}})={\color{green} {\cfrac{6}{9}}}(x-({\color{red}{ -10}}))\implies y+4=\cfrac{6}{9}(x+10)\)

OpenStudy (the_fizicx99):

@jdoe0001 6/9 simplfiys to 2/3

OpenStudy (jdoe0001):

yes :)